# Near-Linear and Parameterized Approximations for Maximum Cliques in Disk Graphs

### v2

@@ 3 Unit Disk Graphs
We describe an algorithm that computes an $(1-\varepsilon)$-approximate clique in a unit disk graph with probability at least $1-\delta$ in $O((n/\varepsilon^{-2})\log^{4}(n)\log(1/\delta))$ expected time. Let $P$ be a set of $n$ points in the plane, and let $\mathcal{D}(P)=\{D(p,1)\mid p\in P\}$ be the set of unit disks whose centers are given by $P$. For convenience, we will sometimes identify disks with their centers. A set $X\subseteq P$ is said to form a clique if all the disks in $\mathcal{D}(X)$ intersect pairwise.
@@
Now, let $C\in\mathcal{C}$. We select two disk centers $p_{1},p_{2}\in P_{C}$ independently and uniformly at random. Let $s$ be the line segment of length 2 that starts at $p_{2}$ and is directed toward $p_{1}$, so that $p_{1}$ lies either on $s$ or on its extension along the same direction. Let $x$ be the other endpoint of $s$, and let $L=D(x,2)\cap D(p_{2},2)$ be the lens with axis $s$; see fig.˜3 for an illustration. We determine the set $P_{L}=L\cap P_{C}$ of disks centers that lie in $L$, and we find an $(1-\varepsilon/2)$-approximate maximum clique $X$ for $P_{L}$, using theorem˜4 from section˜2. The lens $L$ is partitioned by segment $s$ into two parts: the part to the left of $s$ and the part to the right of $s$. Without loss of generality, the points on $s$ belong to the left part. Any two points in the same part have a distance at most $2$. Thus, all disks with centers in the same part form a clique. In other words, the intersection graph of unit disks centered at $P_{L}$ is a co-bipartite graph. Consequently, as we show below, the clique $X$ is an $(1-\varepsilon)$-approximate clique for $P_{C}$ with probability $\Omega(\varepsilon)$.
@@
Let $X^{*}$ be a maximum clique in $P_{C}$. Then, with probability $\Omega(\varepsilon)$, we have $|X|\geq(1-\varepsilon))\,|X^{*}|.$
@@
First, by ˜6, we know that $|X^{*}|\geq|P_{C}|/25$. Thus, with probability at least $1/25$, we have that $p_{1}\in X^{*}$. Now, we sort the centers in $X^{*}$ in the increasing order of their distance from $p_{1}$, and let $\pi^{*}$ be the resulting ordering of $X^{*}$. Let $X^{+}$ be those centers from $X^{*}$ that have rank at least $(1-\varepsilon/2)|X^{*}|$ in $\pi^{*}$, and let $X^{-}=X^{*}\setminus X^{+}$ (in other words, $X^{+}$ contains the $(\varepsilon/2)|X^{*}|$ points in $X^{*}$ that are furthest away from $p_{1}$, and $X^{-}$ contains the $(1-\varepsilon/2)|X^{*}|$ remaining points). Since
@@
As $p_{2}\in X^{+}$, and since all the points in $X^{+}$ are at least as far away from $p_{1}$ as all the point in $X^{-}$, it follows that $X^{-}\subseteq D(p_{1},\mathrm{dist}(p_{1},p_{2})\subseteq D(x,2)$. Moreover, since $X^{-}\subseteq X^{*}$ and $p_{2}\in X^{*}$, we also have $X^{-}\subseteq D(p_{2},2)$. This implies that $X^{-}$ lies in the lens $L$ defined by $D(x,2)$ and $D(p_{2},2)$. In particular, $P_{L}$ contains a clique of size $(1-\varepsilon/2)|X^{*}|$, and since $X$ is a $(1-\varepsilon/2)$-approximate clique for $P_{L}$, we have
@@
This means that $X$ is an $\varepsilon$-approximate maximum clique for $P_{C}$, as claimed. ∎
@@
Let $P$ be a set of $n$ points in the plane, let $\varepsilon>0$ be a parameter. There is a randomized algorithm that runs in expected time $O((n/\varepsilon^{2})\log^{4}n)$ and computes an $(1-\varepsilon)$-approximate clique for $\mathcal{D}(P)$ with probability at least $1/2$.
@@
Let $C\in\mathcal{C}$ be a nonempty grid cell. By Lemma˜7, and using Theorem˜4, one can compute, in $O((|P_{C}|/\varepsilon)\log^{4}n)$ expected time, a clique for $P_{C}$ that, with probability $\Omega(\varepsilon)$, is an $(1-\varepsilon)$-approximate clique for $P_{C}$. By repeating this process $O(1/\varepsilon)$ times, and selecting the largest of the resulting cliques, one can compute in $O((|P_{C}|/\varepsilon^{2})\log^{4}n)$ expected time, a clique that is an $(1-\varepsilon)$-approximate clique for $P_{C}$ with probability at least $1/2$.
We do this for $P_{C}$ of every nonempty cell $C\in\mathcal{C}$, and we return the largest of the resulting cliques. (Note that all non-empty cells can be easily found in $O(n\log n)$ time [Wang23]). By ˜5, there exists a cell $C\in\mathcal{C}$ such that $P_{C}$ contains a max-clique of $P$, so with probability at least $1/2$, this gives an $(1-\varepsilon)$-approximate maximum clique for $P$. The total running time is proportional to
$\sum_{C\in\mathcal{C}}\frac{|P_{C}|}{\varepsilon^{2}}\log^{4}n=\frac{\sum_{C\in\mathcal{C}}|P_{C}|}{\varepsilon^{2}}\log^{4}n=O((n/\varepsilon^{2})\log^{4}n),$
@@
By a standard argument, for any value $\delta>0$, we can raise the probability of success to $1-\delta$ by independently repeating the algorithm of Theorem˜8 $O(\log(1/\delta))$ times and by taking the best solution. This increases the running time by an $O(\log(1/\delta))$ factor.
@@ 4 Disk graphs with different radii
We first review the algorithm by Keil and Mondal [keil_et_al:LIPIcs.SoCG.2025.63]. The main idea is simple and elegant, and it goes as follows. Suppose that there are $t$ distinct radii $r_{1},\dots,r_{t}$, in increasing order. One can enumerate all radii that appear in a maximum clique and, for each radius $r_{i}$, enumerate all disk pairs to find the one with centers $a_{i},b_{i}$ that are, respectively, leftmost and rightmost (in the optimum) along the $x$-axis. Let $X$ be the set of these $2t$ disks. For each pair of centers $a_{i},b_{i}$, consider their vertical slab and find the set $L_{i}$ of all radius-$r_{i}$ disks that intersect all disks in $X$ and have their centers within the slab and above the line segment $a_{i}b_{i}$. Similarly, find the set $R_{i}$ of all radius-$r_{i}$ disks that intersect all disks in $X$ and have their centers in the slab and below the line segment $a_{i}b_{i}$. Take $L=L_{1}\cup L_{2}\cup\dots\cup L_{t}$ and $R=R_{1}\cup R_{2}\cup\dots\cup R_{t}$. A simple geometric argument shows that the graph induced by $L\cup R$ is co-bipartite, and thus, similar to the case of unit disks, one can find a maximum clique by finding a maximum matching. The algorithm runs in time $O(2^{t}n^{2t}(f(n)+n^{2}))$, where $f(n)$ is the time to compute a maximum matching in an $n$-vertex bipartite graph. The factor $n^{2t}$ comes from the enumeration to find the leftmost and rightmost disks in the optimal solution for every radius $r_{i}$.
In this section, we give an efficient parameterized approximation scheme w.r.t. the number $t$ of different radii and $\varepsilon$. There are two places where we can speed up the algorithm described above if approximate solutions are allowed. First, instead of finding the leftmost and rightmost disks for each radius by enumeration, we use random sampling to find two centers $a_{i},b_{i}$ that are close (in terms of rank) to the leftmost and rightmost disks in the optimal clique. Second, as in the case of unit disk graphs, we can replace finding a maximum clique by finding an approximate maximum clique using Theorem˜4.
@@ 4.1 A special case: Disks of every radius contribute to the maximum clique
    For each radius $r_{i}$: uniformly at random pick two disks $D(a_{i},r_{i})$ and $D(b_{i},r_{i})$ from $\mathcal{D}^{\prime}_{i}=\mathcal{D}_{i}\cap\mathcal{D}^{\prime}$. Let $X$ denote the selected disks. Without loss of generality, assume that $a_{i}$ is to the left of $b_{i}$. We interpret $a_{i}$ as the leftmost disk of radius $r_{i}$ in the clique, and $b_{i}$ as the rightmost disk in the clique.
@@
    Use the algorithm from Theorem 4 to compute an $(1-\varepsilon/2)$-approximate clique in $G(L\cup R)$. Note that this is possible, as $G(L\cup R)$ is co-bipartite, see Lemma˜10 below.
@@
The proof that $L$ and $R$ form a clique respectively follows directly from [keil_et_al:LIPIcs.SoCG.2025.63, Lemma 3.1]. Next, all disks in $D^{*}$ intersect $D(o,r_{1})$, and $D(a_{i},r_{i}),D(b_{i},r_{i})$, for all $i$. Thus the disks of radius $r_{i}$ of $D^{*}$ with the conditions specified in the lemma meet the requirements for $L_{i}$ or $R_{i}$ and thus are included in $L\cup R$. ∎
@@
For each disk $D(o,r_{1})$, finding the disks that intersect it can be done easily in $O(n)$ time. But if we repeat this step $m_{1}=\Theta(n/k\varepsilon)$ times, this will be too costly. Instead, we pre-process the disks into a data structure and answer a 2D circular range query, which returns the points within a query disk. In particular, for each radius $r_{i}$ we process the centers of the disks in $\mathcal{D}_{i}$ into a size $O(|\mathcal{D}_{i}|)$ data structure using the construction in [Afshani2009-ih] such that one can report all centers of the disks in $\mathcal{D}_{i}$ within distance $r_{1}+r_{i}$ from $o$ in $O(\log n+\ell_{i})$ time, where $\ell_{i}$ is the size of the output. The construction time for the data structure for $\mathcal{D}_{i}$ is $O(|\mathcal{D}_{i}|\log|\mathcal{D}_{i}|)$. Thus total construction time for all $t$ different radii is $O(n\log n)$. We answer $tm_{1}$ queries and the total query time is $O(tm_{1}\log n+m_{1}k^{*})$.
@@ 4.2 The complete algorithm
Let $\mathcal{D}$ be a set of $n$ disks, let $t$ be the number of distinct radii in $\mathcal{D}$, and let $\varepsilon\in 0$ be a parameter. With a constant success probability, we can compute a $(1-\varepsilon)$-approximate maximum clique in $G(\mathcal{D})$ in expected time $\tilde{O}\left((\frac{t}{\varepsilon})^{O(t)}n\right)$.
@@
Consider the optimal solution $D^{*}$ with $k^{*}$ disks, we remove all the disks of radius $r_{i}$ from $D^{*}$ if there are at most $\varepsilon k^{*}/(2t)$ of them in $D^{*}$. This leaves a clique $\bar{D}\subseteq D^{*}$ with $(1-\varepsilon/2)k^{*}\leq|\bar{D}|\leq k^{*}$. $\bar{D}$ has at least $\varepsilon k^{*}/2t\geq\varepsilon k/2t$ of disks in radius $r_{i}$ if $r_{i}$ appears at all. We now aim to find a $1-\varepsilon/2$ approximation to $\bar{D}$, which is a $1-\varepsilon$ approximation to $D^{*}$. Of course, we do not know in advance what those radii are that appear less than $\varepsilon k^{*}/(2t)$ in $D^{*}$, which is some (fixed but unknown) optimal solution, and therefore we will brute-force through all subsets of radii, at the cost of an extra factor of $2^{t}$ in the running time.
@@
The running time comes from repeating Lemma 11 with parameters $\varepsilon/2$ and $k/t$, for $2^{t}$ times. Notice that the construction of the circular range query data structure is only done once at the initialization step. But all other steps will be repeated $2^{t}$ times. Further, $k^{*}/2\leq k\leq k^{*}$. The total running time is obtained from applying Lemma 11.

### v3

@@ 3 Unit Disk Graphs
We describe an algorithm that computes an $(1-\varepsilon)$-approximate clique in a unit disk graph with probability at least $1-\delta$ in $O((n/\varepsilon^{2})\log(n)\log(1/\delta))$ expected time. Let $P$ be a set of $n$ points in the plane, and let $\mathcal{D}(P)=\{D(p,1)\mid p\in P\}$ be the set of unit disks whose centers are given by $P$. For convenience, we will sometimes identify disks with their centers. A set $X\subseteq P$ is said to form a clique if all the disks in $\mathcal{D}(X)$ intersect pairwise.
@@
Now, let $C\in\mathcal{C}$. We select two disk centers $p_{1},p_{2}\in P_{C}$ independently and uniformly at random. Let $s$ be the line segment of length 2 that starts at $p_{2}$ and is directed toward $p_{1}$, so that $p_{1}$ lies either on $s$ or on its extension along the same direction. Let $x$ be the other endpoint of $s$, and let $L=D(x,2)\cap D(p_{2},2)$ be the lens with axis $s$; see fig.˜3 for an illustration. We determine the set $P_{L}=L\cap P_{C}$ of disks centers that lie in $L$, and we find an $(1-\varepsilon/2)$-approximate maximum clique $X$ for $P_{L}$, using theorem˜6 from section˜2. The lens $L$ is partitioned by segment $s$ into two parts: the part to the left of $s$ and the part to the right of $s$. Without loss of generality, the points on $s$ belong to the left part. Any two points in the same part have a distance at most $2$. Thus, all disks with centers in the same part form a clique. In other words, the intersection graph of unit disks centered at $P_{L}$ is a co-bipartite graph. Consequently, as we show below, the clique $X$ is an $(1-\varepsilon)$-approximate clique for $P_{C}$ with probability $\Omega(\varepsilon)$.
@@
Let $X^{*}$ be a maximum clique in $P_{C}$. Then, with probability $\Omega(\varepsilon)$, we have $|X|\geq(1-\varepsilon)|X^{*}|$.
@@
First, by ˜8, we know that $|X^{*}|\geq|P_{C}|/25$. Thus, with probability at least $1/25$, we have that $p_{1}\in X^{*}$. Now, we sort the centers in $X^{*}$ in the increasing order of their distances from $p_{1}$, and let $\pi^{*}$ be the resulting ordering of $X^{*}$. Let $X^{+}$ be those centers from $X^{*}$ that have rank at least $(1-\varepsilon/2)|X^{*}|$ in $\pi^{*}$, and let $X^{-}=X^{*}\setminus X^{+}$ (in other words, $X^{+}$ contains the $(\varepsilon/2)|X^{*}|$ points in $X^{*}$ that are furthest away from $p_{1}$, and $X^{-}$ contains the $(1-\varepsilon/2)|X^{*}|$ remaining points). Since
@@
As $p_{2}\in X^{+}$, and since all the points in $X^{+}$ are at least as far away from $p_{1}$ as all the points in $X^{-}$, it follows that $X^{-}\subseteq D(p_{1},\mathrm{dist}(p_{1},p_{2}))\subseteq D(x,2)$. Moreover, since $X^{-}\subseteq X^{*}$ and $p_{2}\in X^{*}$, we also have $X^{-}\subseteq D(p_{2},2)$. This implies that $X^{-}$ lies in the lens $L$ defined by $D(x,2)$ and $D(p_{2},2)$. In particular, $P_{L}$ contains a clique of size $(1-\varepsilon/2)|X^{*}|$, and since $X$ is a $(1-\varepsilon/2)$-approximate clique for $P_{L}$, we have
@@
This means that $X$ is an $(1-\varepsilon)$-approximate maximum clique for $P_{C}$, as claimed. ∎
@@
Let $P$ be a set of $n$ points in the plane, let $\varepsilon>0$ be a parameter. There is a randomized algorithm that runs in expected time $O((n/\varepsilon^{2})\log n)$ and computes an $(1-\varepsilon)$-approximate clique for $\mathcal{D}(P)$ with probability at least $1/2$.
@@
Let $C\in\mathcal{C}$ be a nonempty grid cell. By Lemma˜9, and using Theorem˜6, one can compute, in $O((|P_{C}|/\varepsilon)\log n)$ expected time, a clique for $P_{C}$ that, with probability $\Omega(\varepsilon)$, is an $(1-\varepsilon)$-approximate clique for $P_{C}$. By repeating this process $O(1/\varepsilon)$ times, and selecting the largest of the resulting cliques, one can compute in $O((|P_{C}|/\varepsilon^{2})\log n)$ expected time, a clique that is an $(1-\varepsilon)$-approximate clique for $P_{C}$ with probability at least $1/2$.
We do this for $P_{C}$ of every nonempty cell $C\in\mathcal{C}$, and we return the largest of the resulting cliques. (Note that all non-empty cells can be easily found in $O(n\log n)$ time [36]). By ˜7, there exists a cell $C\in\mathcal{C}$ such that $P_{C}$ contains a max-clique of $P$, so with probability at least $1/2$, this gives an $(1-\varepsilon)$-approximate maximum clique for $P$. The total running time is proportional to
$\sum_{C\in\mathcal{C}}\frac{|P_{C}|}{\varepsilon^{2}}\log n=\frac{\sum_{C\in\mathcal{C}}|P_{C}|}{\varepsilon^{2}}\log n=O((n/\varepsilon^{2})\log n),$
@@
By a standard argument, for any value $\delta>0$, we can raise the probability of success to $1-\delta$ by independently repeating the algorithm of Theorem˜10 $O(\log(1/\delta))$ times and by taking the best solution. This increases the running time by an $O(\log(1/\delta))$ factor.
@@ 4 Disk graphs with different radii
We first review the algorithm by Keil and Mondal [29]. The main idea is simple and elegant, and it goes as follows. Suppose that there are $t$ distinct radii $r_{1},\dots,r_{t}$, in increasing order. One can enumerate all radii that appear in a maximum clique and, for each radius $r_{i}$, enumerate all disk pairs to find the one with centers $a_{i},b_{i}$ that are, respectively, leftmost and rightmost (in the optimum) along the $x$-axis. Let $X$ be the set of these $2t$ disks. For each pair of centers $a_{i},b_{i}$, consider their vertical slab and find the set $L_{i}$ of all radius-$r_{i}$ disks that intersect all disks in $X$ and have their centers within the slab and above the line segment $a_{i}b_{i}$. Similarly, find the set $R_{i}$ of all radius-$r_{i}$ disks that intersect all disks in $X$ and have their centers in the slab and below the line segment $a_{i}b_{i}$. Take $L=L_{1}\cup L_{2}\cup\dots\cup L_{t}$ and $R=R_{1}\cup R_{2}\cup\dots\cup R_{t}$. A simple geometric argument shows that the graph induced by $L\cup R$ is co-bipartite, and thus, similar to the case of unit disks, one can find a maximum clique by finding a maximum matching. The algorithm runs in time $O(2^{t}n^{2t}(f(n)+n^{2}))$, where $f(n)$ is the time to compute a maximum matching in an $n$-vertex bipartite graph. The factor $n^{2t}$ comes from the enumeration to find the leftmost and rightmost disks in the optimal solution for every radius $r_{i}$.
In this section, we give an efficient parameterized approximation scheme w.r.t. the number $t$ of different radii and $\varepsilon$. There are two places where we can speed up the algorithm described above if approximate solutions are allowed. First, instead of finding the leftmost and rightmost disks for each radius by enumeration, we use random sampling to find two centers $a_{i},b_{i}$ that are close (in terms of rank) to the leftmost and rightmost disks in the optimal clique. Second, as in the case of unit disk graphs, we can replace finding a maximum clique by finding an approximate maximum clique using Theorem˜6.
@@ 4.1 A special case: Disks of every radius contribute to the maximum clique
    For each radius $r_{i}$: uniformly at random pick two disks $D(a_{i},r_{i})$ and $D(b_{i},r_{i})$ from $\mathcal{D}^{\prime}_{i}=\mathcal{D}_{i}\cap\mathcal{D}^{\prime}$. Let $X$ denote the set of the selected disks for all $r_{i}$. Without loss of generality, assume that $a_{i}$ is to the left of $b_{i}$. We interpret $a_{i}$ as the leftmost disk of radius $r_{i}$ in the clique, and $b_{i}$ as the rightmost disk in the clique.
@@
    Use the algorithm from Theorem 6 to compute an $(1-\varepsilon/2)$-approximate clique in $G(L\cup R)$. Note that this is possible, as $G(L\cup R)$ is co-bipartite, see Lemma˜12 below.
@@
The proof that $L$ and $R$ form a clique respectively follows directly from [29, Lemma 3.1]. Next, all disks in $D^{*}$ intersect $D(o,r_{1})$, and $D(a_{i},r_{i}),D(b_{i},r_{i})$, for all $i$. Thus the disks of radius $r_{i}$ of $D^{*}$ with the conditions specified in the lemma meet the requirements for $L_{i}$ or $R_{i}$ and thus are included in $L\cup R$. ∎
@@
For each disk $D(o,r_{1})$, finding the disks that intersect it can be done easily in $O(n)$ time. But if we repeat this step $m_{1}=\Theta(n/k\varepsilon)$ times, this will be too costly. Instead, we pre-process the disks into a data structure and answer a 2D circular range query, which returns the points within a query disk. In particular, for each radius $r_{i}$ we process the centers of the disks in $\mathcal{D}_{i}$ into a size $O(|\mathcal{D}_{i}|)$ data structure using the construction in [1] such that one can report all centers of the disks in $\mathcal{D}_{i}$ within distance $r_{1}+r_{i}$ from $o$ in $O(\log n+\ell_{i})$ time, where $\ell_{i}$ is the size of the output. The construction time for the data structure for $\mathcal{D}_{i}$ is $O(|\mathcal{D}_{i}|\log|\mathcal{D}_{i}|)$. Thus total construction time for all $t$ different radii is $O(n\log n)$. We answer $tm_{1}$ queries and the total query time is $O(tm_{1}\log n+m_{1}k^{*})$.
@@ 4.2 The complete algorithm
Let $\mathcal{D}$ be a set of $n$ disks, let $t$ be the number of distinct radii in $\mathcal{D}$, and let $\varepsilon\in(0,1)$ be a parameter. With a constant success probability, we can compute a $(1-\varepsilon)$-approximate maximum clique in $G(\mathcal{D})$ in expected time $\tilde{O}\left((\frac{t}{\varepsilon})^{O(t)}n\right)$.
@@
Consider the optimal solution $D^{*}$ with $k^{*}$ disks and assume that all disks of radius $r_{i}$ are removed from $D^{*}$ if there are at most $\varepsilon k^{*}/(2t)$ of them in it. This leaves a clique $\bar{D}\subseteq D^{*}$ with $(1-\varepsilon/2)k^{*}\leq|\bar{D}|\leq k^{*}$. Note that $\bar{D}$ has at least $\varepsilon k^{*}/2t\geq\varepsilon k/2t$ of disks in radius $r_{i}$, if $r_{i}$ appears at all. We now aim to find a $1-\varepsilon/2$ approximation to $\bar{D}$, which is a $1-\varepsilon$ approximation to $D^{*}$. Of course, we do not know in advance what those radii are that appear less than $\varepsilon k^{*}/(2t)$ in $D^{*}$, which is some (fixed but unknown) optimal solution, and therefore we will brute-force through all subsets of radii, at the cost of an extra factor of $2^{t}$ in the running time.
@@
The running time comes from repeating Lemma 13 with parameters $\varepsilon/2$ and $k/t$, for $2^{t}$ times. Notice that the construction of the circular range query data structure is only done once at the initialization step. But all other steps will be repeated $2^{t}$ times. Further, $k^{*}/5\leq k\leq k^{*}$. The total running time is obtained from applying Lemma 13.